OpenStudy (anonymous):

SOMEBODY HELP ME PLZ Find f(x) and g(x) so that the function can be described as y = f(g(x)). y=9/sqrt{5x+5}

OpenStudy (anonymous):

I always get these wrong every time I try. D: My choices are: f(x) = 9/squarerootx., g(x) = 5x + 5 f(x) = squareroot of 5x+5, g(x) = 9 f(x) = 9/x, g(x) = 5x + 5 f(x) = 9, g(x) = squareroot x+5